OpenStudy (anonymous):

Find the Product:

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 }x ^{3}(\frac{ -7 }{ 2 }x ^{2})\]

OpenStudy (jdoe0001):

hmmm actuall is a 3

OpenStudy (jdoe0001):

\(\bf \cfrac{ 4 }{ 5 }x ^{3}\left(\cfrac{ -7 }{ 2 }x ^{2}\right)\implies \cfrac{ 4x^3}{ 5 }\cdot \cfrac{ -7 x ^{2}}{ 2 }\implies \cfrac{4x^3\cdot -7 x ^{2}}{ 5\cdot 2 }\implies \cfrac{4\cdot -7\cdot x^3\cdot x ^{2}}{ 5\cdot 2 }\)

OpenStudy (anonymous):

I thought you multiplied the feactions and then addded the exponents?